# Quantum hash function using discrete-time quantum walk on Hanoi network

### v1

@@ II Quantum walk on Hanoi network
Hanoi network (HN) is generated from the sequence of numbered disks of the Tower-of-Hanoi problem [40]. Based on the number of edges at each vertex, there exist two important type of networks [37]: Hanoi network of degree three (HN3) and of degree four (HN4). Extensions of these typical Hanoi networks with even more degrees exist, but in this article, we only consider HN4 network for our study of the hash function. The main structure of this network is a one-dimensional periodic lattice of $N_{v}=2^{n}$ vertices. Like the one-dimensional lattice, it has two regular edges at each vertex $1\leq x_{v}\leq 2^{n}$, which are connected to its two nearest neighbor vertices. Additionally, it also has long-range edges, which give the small-world structure to the Hanoi network.
@@
One of the important properties of the quantum walk on a finite one-dimensional lattice with $t$ time steps is that, for even $N$, probability distribution at location $x$ is zero [15] when $x+t$ is odd as can be seen from fig. 2(a). Therefore, the probability distribution is zero at even or add vertices for odd or even $t$ respectively. However, for odd $N$ the probability distribution is nonzero at all vertex locations for large enough $t$. So, normally, one cannot use even length one-dimensional periodic lattice for the generation of hash function, because it leads to very high collision rates due the predictable periodic zero probability distributions. The HN4 network, which we use in this study, allows us to consider even-length one-dimensional periodic lattice with no periodic zero-probability distribution as can be seen from fig. 2(b). Also, because of its two long-range edges we can manipulate probability distribution through the coin operator in such a way that the generated hash function has very high collision resistance. Note that, lively quantum walk with even lively-parameter studied in ref. [31] also does not have periodic zero probability distributions.
@@ III Hash function with quantum walk on Hanoi network
For example, for a message $m=`110100^{\prime}$, the final state would be $|\psi_{f}\rangle=\mathcal{U}_{0}\mathcal{U}_{0}\mathcal{U}_{1}\mathcal{U}_{0}\mathcal{U}_{1}\mathcal{U}_{1}|\psi_{in}\rangle$. The iteration steps for the generation of our hash function, obtained from eq. (10) increases linearly with the bit-length $s$ of the message. Although, the results presented in this article are based on noiseless analysis, in real-life scenarios, a long bit-length can transform a quantum probability distribution of the quantum walk into a classical Gaussian-like distribution in the presence of unitary noise [41]. Correction methods such as error correction [42], and decoherence free sub-spaces [43] require significant overhead, which is a luxury in the current NISQ era. The effect of noise can be minimized by reducing the number of iteration steps of our scheme significantly by adopting an approach from ref. [33], where each evolution operator is controlled by two bits of the message instead of one in eq. (10), without compromising the performance our hash function. The final state for the hash function then becomes
@@
where $\mathcal{U}_{ij}=S_{i}\left(C_{ij}\otimes I\right)$. Four different coin operators $C_{00},C_{01},C_{10},C_{11}$, needed for this method, can be obtained by appropriately choosing two parameters associated with the long-range edges in the underlying coin state. For example, for a message $m=`110100^{\prime}$, the final state would now be $|\psi_{f}\rangle=\mathcal{U}_{00}\mathcal{U}_{10}\mathcal{U}_{11}|\psi_{in}\rangle$. The required iteration steps for the final state $|\psi_{f}\rangle$ can further be reduced by adopting other similar methods presented in ref. [33].
@@ IV Performance of hash function
In general, our proposed hash function has a bit-length $L=N_{v}\times k$, where $N_{v}$ is the number of vertices $x_{v}$ of the HN4 network and $k$ is the bit-length of each probability amplitude $h_{x_{v}}$. There is no conclusive analytical proof for the security of the hash function. In the following sub-sections we perform several standard statistical analysis [24, 25, 26, 29, 30, 31, 32, 33, 34, 35], which show that the hash function is secured and highly collision resistant.
@@ IV.2 Diffusion and confusion properties
Figure 5: Experimental (green) and theoretical (red) collision rate ($\%$) based on $N=10000$ pair of randomly chosen $1.5N_{v}$-bit length messages as function of the number of vertices of the HN4 network for $l=7,k=16$.
@@ IV.4 Collision analysis
$W^{T}(\omega)$ and $W^{E}(\omega)$ presented in table 4 shows that the experimental values are close to the theoretical values. The experimental collision rare of $0.05\%$ is very close to the theoretical collision rate of $0.02\%$, indicating very strong collision resistance. Note that such high collision resistance of our hash function is attributed to the suitable choice of values for the long-range-edge weights $l_{00},\tilde{l}_{00},l_{11},\tilde{l}_{11}$. The set of values $l_{00}=0.01,\tilde{l}_{00}=1.0,l_{11}=0.1,\tilde{l}_{11}=0.01$ chosen heuristically in this article is one of the many such suitable set of values for which the experimental collision rate $W^{E}(\omega)$ is close to the theoretical collision rate $W^{T}(\omega)=0.02\%$. It is also possible to obtain optimum values for the long-range-edge weights by optimizing the collision rate of the hash scheme with respect to the corresponding theoretical collision rate.
@@ IV.5 Resistance to birthday attack
In this paper we have discussed hash values with a bit-length of $L=N_{v}\times k=16\times 16=256$, which means we need $2^{L/2}=2^{128}$ trials to find if two hash values collide with a probability of $0.5$, which suggests high resistance to birthday attack. Moreover, we can further increase bit-length of the hash value by increasing the length of the lattice $N_{v}$. In cryptography, fixed-size classical hash functions up-to $512$-bit hash values are currently being used in practice. Our hash function can easily be extended to $512$-bit hash value, corresponding to $N_{v}=32$, and beyond with high collision resistance property by appropriately tuning the precision level. Fig. 5 shows it has good scaling property as the collision rate of the hash function, represented by the dashed green curve is very close to the theoretical collision rate, represented by the solid red curve.
@@ V Comparison with other quantum walk based hash schemes
One of the most important indicators for a secured hash function is the collision rate. The lower the collision rate the better is the hash function. A comparison of the collision rates(experimental and theoretical) of different hash schemes is presented in table 5, where the theoretical values are obtained using the binomial distribution in eq. (16). Since the required details for calculating $W^{T}$ for ref. [24] is not available, we calculate its theoretical value of collision in table 5, assuming $N_{v}=25$ vertices and each vertex contributing $8$ bit to a $200$-bit hash value and performing collision test for $N=10000$ times. Collision rate of our hash scheme, presented in the seventh column of table 5, shows that the experimental value $0.05$ is very close to the theoretical value $0.02$. Collision rate of our scheme can further be improved using higher precision label as shown in fig. 5. A comparison of the resource requirements of our scheme along with others are presented in table 6. Two important aspects of our proposed scheme, which are responsible for such high collision resistance property, are the following: Firstly, we have introduced two additional long-range edges to the two regular edges that allows us to tune the parameters in such a way that the probability distribution can be modified. Secondly, in addition to the coin operator we also control the shift operator by the message bits as opposed to the other quantum walk based hash functions which only control the coin operator by the message.

### v2

@@ II Quantum walk on Hanoi network
Hanoi network (HN) is generated from the sequence of numbered disks of the Tower-of-Hanoi problem [40]. Based on the number of edges at each vertex, there exist two important type of networks [37]: Hanoi network of degree three (HN3) and of degree four (HN4). Extensions of these typical Hanoi networks with even more degrees exist, but, only HN3 and HN4 networks are preferable, since they require the smallest number of qubits ($\mbox{qubits}=2$) to represent the edges of the one-dimensional periodic lattice with extra long-range edges, which is preferable for the implementation in NISQ devices that have limited and qubits and low connectivity among qubits. In this article we have considered the HN4 network because the coin operator can be easily implemented in quantum circuits since all the four basis states of the two-qubit coin space are utilized, however we have numerically checked that the quantum hash function using the HN3 network also has similar performance. The main structure of this network is a one-dimensional periodic lattice of $N_{v}=2^{n}$ vertices. Like the one-dimensional lattice, it has two regular edges at each vertex $1\leq x_{v}\leq 2^{n}$, which are connected to its two nearest neighbor vertices. Additionally, it also has long-range edges, which give the small-world structure to the Hanoi network.
@@
One of the important properties of the quantum walk on a finite one-dimensional lattice with $t$ time steps is that, for even $N$, the probability distribution at location $x$ is zero [15] when $x+t$ is odd, as can be seen from Fig. 2(a). Therefore, the probability distribution is zero at even or odd vertices for odd or even $t$ respectively. However, for odd $N$ the probability distribution is nonzero at all vertex locations for large enough $t$. So, normally, one cannot use even-length one-dimensional periodic lattice [31] for the generation of a hash function, because it leads to very high collision rates due to the predictable periodic zero probability distributions. The HN4 network, which we use in this study, allows us to consider an even-length one-dimensional periodic lattice with no periodic zero-probability distribution as can be seen from Fig. 2(b). Also, because of its two long-range edges we can manipulate probability distribution through the coin operator in such a way that the generated hash function has very high collision resistance. Note that, lively quantum walk with even lively-parameter studied in ref. [31] also does not have periodic zero probability distributions.
@@ III Hash function with quantum walk on Hanoi network
For example, for a message $m=`110100^{\prime}$, the final state would be $|\psi_{m}\rangle=\mathcal{U}_{0}\mathcal{U}_{0}\mathcal{U}_{1}\mathcal{U}_{0}\mathcal{U}_{1}\mathcal{U}_{1}|\psi_{in}\rangle$. The iteration steps for the generation of our hash function, obtained from eq. (10) increases linearly with the bit-length $s$ of the message.
Although, the results presented in this article are based on noiseless analysis, in real-life scenarios, a longer bit-length, corresponding to more iteration steps, can transform a quantum probability distribution of the quantum walk into a classical Gaussian-like distribution in the presence of unitary noise [41]. Noise associated with more iterations steps significantly degrades the probability amplitudes in Grover search algorithms also, which can be mitigated with strategies involving less iteration steps [42]. Correction methods such as error correction [43], and decoherence free sub-spaces [44] require significant overhead, which is a luxury in the current NISQ era. The effect of noise can be minimized by reducing the number of iteration steps of our scheme significantly by adopting an approach from ref. [33], where each evolution operator is controlled by two bits of the message instead of one in eq. (10), without compromising the performance of our hash function. The final state for the hash function then becomes
@@
where $\mathcal{U}_{ij}=S_{i}\left(C_{ij}\otimes I\right)$. Four different coin operators $C_{00},C_{01},C_{10},C_{11}$, needed for this method, can be obtained by appropriately choosing two parameters associated with the long-range edges in the underlying coin state. For example, for a message $m=`110100^{\prime}$, the final state would now be $|\psi_{f}\rangle=\mathcal{U}_{00}\mathcal{U}_{10}\mathcal{U}_{11}|\psi_{in}\rangle$. The required iteration steps for the final state $|\psi_{f}\rangle$ can further be reduced by adopting other similar methods presented in ref. [33]. We acknowledge that a comprehensive performance validation under realistic noisy conditions remains an essential issue. Future work will involve detailed simulations to assess the impact of varying noise levels on collision rates, distribution uniformity, and the effectiveness of our proposed scheme.
@@ IV Performance of hash function
In general, our proposed hash function has a bit-length $L=N_{v}\times k$, where $N_{v}$ is the number of vertices $x_{v}$ of the HN4 network and $k$ is the bit-length of each probability amplitude $h_{x_{v}}$. There is no conclusive analytical proof for the security of the hash function. However, the process of our quantum walk based hash function involves measurement of final quantum state in computational basis and in the post-processing stage module $2^{k}$ operation, both of which being many-to-one mappings, are irreversible one-way processes. Therefore, it is resistant to preimage attack, as given a hash value $y$ it is hard to find its corresponding message $m$, such that $H(m)=y$. It is also resistant to second preimage attacks, since given a message $m_{1}$ it is computationally infeasible because of the irreversible nature of the involved processes, to find another message $m_{1}$ such that $H(m_{1})=H(m_{2})$.
In the following sub-sections we perform several standard statistical analysis [24, 25, 26, 29, 30, 31, 32, 33, 34, 35], which show that the hash function is secured and highly collision resistant.
@@ IV.2 Diffusion and confusion properties
Figure 5: Experimental (green) and theoretical (red) collision rate based on $N=10000$ pair of randomly chosen $1.5N_{v}$-bit length messages as function of the number of vertices of the HN4 network for $l=7,k=16$.
@@ IV.3 Uniform distribution
Figure 6: Experimental (solid green) and theoretical (dashed green) collision rates as function of $N=10000$ pair of randomly chosen messages of varying bit-length with HN4 network for $l=6,k=16$.
@@
Figure 7: Comparison of collision rates for five randomly chosen set of long-range weights $\lambda_{00},\tilde{\lambda}_{00},\lambda_{11},\tilde{\lambda}_{11},$ of the HN4 network.
@@ IV.4 Collision analysis
$W^{T}(\omega)$ and $W^{E}(\omega)$ presented in table 4 shows that the experimental values are close to the theoretical values. The experimental collision rare of $c_{E}=W^{E}(\omega)\times 100=0.05\%$ is very close to the theoretical collision rate of $c_{T}=W^{T}(\omega)\times 100=0.02\%$, indicating very strong collision resistance. Note that such high collision resistance of our hash function is attributed to the suitable choice of values for the long-range-edge weights $l_{00},\tilde{l}_{00},l_{11},\tilde{l}_{11}$. The set of values $l_{00}=0.01,\tilde{l}_{00}=1.0,l_{11}=0.1,\tilde{l}_{11}=0.01$ chosen heuristically in this article is one of the many such suitable set of values for which the experimental collision rate $c_{E}$ is close to the theoretical collision rate $c_{T}=0.02\%$. The only requirement is that these parameters cannot be zero, since for zero long-range weights HN4 network reduces to a cycle, which has $100\%$ collision rate. In Fig. 7 we evaluated collision rates for five randomly chosen sets of values of the normalized long-range weights $(\lambda_{00}=\mathcal{N}_{0}\sqrt{l_{00}},\tilde{\lambda}_{00}=\mathcal{N}_{0}\sqrt{\tilde{l}_{00}},\lambda_{11}=\mathcal{N}_{1}\sqrt{l_{11}},\tilde{\lambda}_{11}=\mathcal{N}_{1}\sqrt{\tilde{l}_{11}})$ and found that the collision rates are very close to the theoretical value $c_{T}=0.02$ despite wide variations in the parameters, implying any random choices of the long-range weights are suitable for the hash function. We also evaluated trace distance $D_{m1,m2}=\sqrt{1-|\langle\psi_{m1}|\psi_{m2}\rangle|^{2}}$ between the quantum output states $|\psi_{m1}\rangle,|\psi_{m2}\rangle$ corresponding to $N=10000$ randomly chosen pair of messages $(m1,m2)$ and found that trace distance satisfies $D_{m1,m2}\geq 1-\epsilon$ with $\epsilon\leq 0.25$, suggesting strong separability between the output quantum states.
@@ IV.5 Resistance to birthday attack
In this paper we have discussed hash values with a bit-length of $L=N_{v}\times k=16\times 16=256$, which means we need $2^{L/2}=2^{128}$ trials to find if two hash values collide with a probability of $0.5$, which suggests high resistance to birthday attack. Moreover, we can further increase bit-length of the hash value by increasing the length of the lattice $N_{v}$. In cryptography, fixed-size classical hash functions up-to $512$-bit hash values are currently being used in practice. Our hash function can easily be extended to $512$-bit hash value, corresponding to $N_{v}=32$, and beyond with high collision resistance property by appropriately tuning the precision level as can be seen from Fig. 5. It has good scaling property as the collision rate of the hash function, represented by the dashed green curve is very close to the theoretical collision rate, represented by the solid red curve. The parameters $l_{00},\tilde{l}_{00},l_{11},\tilde{l}_{11}$ remain same as the previous experiments in the analysis of Fig. 5, while the numbers of vertices $N_{v}=8,16,32,64,128,256,512$ are considered for the evaluation.
Figure 8: Comparison of collision rates for five randomly chosen set of amplitudes $a_{1},a_{2},a_{3},a_{4}$ of the initial coin state.
@@ V Comparison with other quantum walk based hash schemes
One of the most important indicators for a secured hash function is the collision rate. The lower the collision rate the better is the hash function. A comparison of the collision rates(experimental and theoretical) of different hash schemes is presented in table 5, where the theoretical values are obtained using the binomial distribution in eq. (16). Since the required details for calculating $W^{T}$ for ref. [24] is not available, we calculate its theoretical value of collision in table 5, assuming $N_{v}=25$ vertices and each vertex contributing $8$ bit to a $200$-bit hash value and performing collision test for $N=10000$ times. Collision rate of our hash scheme, presented in the seventh column of table 5, shows that the experimental value $0.05$ is very close to the theoretical value $0.02$. Collision rate of our scheme can further be improved using higher precision label as shown in fig. 5.
A comparison of the resource requirements of our scheme along with others are presented in table 6.
Another advantage of our scheme is that, it works well for short messages also, as demonstrated in Fig. 6. The experimental collision rate, represented by the dashed green curve for $N_{v}=16$ and message bit-length 4204-20, remains close to the theoretical collision rate represented by the solid green line. Although, for very short message bit-length <4<4, the collision rate increases significantly, because the probability does not have enough time to spread. On the other hand since there is predictable zero probability amplitudes for quantum walk on even-length one-dimensional lattice as already discussed in Section II, which results in $100\%$ collision rate for the hash function. This is the reason most of the quantum walk based hash schemes based on one-dimensional periodic lattices cannot use even-length lattice at all. Also they cannot use message bit-length $N_{v}$ for odd-length one-dimensional periodic lattice as the collision rate is very high.
Although the initial state for the analysis is chosen to be equal superposition of all the coin basis states at the origin of the vertex, i.e., $|\psi_{in}\rangle=\frac{1}{2}\sum_{i=0}^{3}|i\rangle\otimes|x_{v}=0\rangle$ for simplicity, however, we can randomly choose the amplitudes $a_{1},a_{2},a_{3},a_{4}$ of the initial state $|\psi_{in}\rangle=\left(a_{1}|0\rangle+a_{2}|1\rangle+a_{3}|2\rangle+a_{4}|3\rangle\right)\otimes|x_{v}=0\rangle$ to carry out our analysis without compromising performance. In Fig. 8 we evaluated collision rates for five randomly chosen sets of amplitudes and found that the collision rates are very close to the theoretical value $c_{T}=0.02$ despite wide variations in the amplitudes, implying any random choices of the initial state are suitable for the hash function.
Two important aspects of our proposed scheme, which are responsible for such high collision resistance property, are the following: Firstly, we have introduced two additional long-range edges to the two regular edges that allow us to tune the parameters in such a way that the probability distribution can be modified. Secondly, in addition to the coin operator we also control the shift operator by the message bits as opposed to the other quantum walk based hash functions which only control the coin operator by the message.